OpenStudy (anonymous):

Whats the answer to (-2/3)(-7/8)(-4/5)?

OpenStudy (anonymous):

do you know how to multiply?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

multiplying all numerators and denominators separately -2*-7*-4=-56 3*8*5=120 answer is -56/120=-7/15